40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 282     Entry time: Mon Jan 30 15:16:54 2017
Author: Gabriele 
Type: General 
Category: Daily Progress 
Subject: Four samples in chamber 
  • checked that the high voltage paths are working:
    • the DAC signal is properly received and amplified by the Trek 2220 HV amplifier
    • the HV switch box is working as expected: I can switch on and off each path individiually using the DAC signals
  • installed four samples (refer to the optical drawing for the numbering)
    • S1600472 in bay 1
    • S1600478 in bay 2
    • S1600473 in bay 3
    • S1600480 in bay 4

  • balanced all samples and centered all optical levers. There is a problem with the fifth connector in the ADC interface board (or maybe with one ADC channel). So one of the quadrants of QPD4 was not working. I moved the connector to number 6 and updated the model. Everything looks fine now.
  • the power in the four QPD is quite diverse: this is not unexpected, since it depends on the interference of the reflection from the two surfaces of the disk

  • installed the picomotor controllers on the bench, connected them to the net using the switch on the top of the clean room. For the moment being they are using DHCP.
  • connected all picomotors to the controllers.
    • Bottom controller, from left to right: QPD1x, QPD1y, QPD2x, QPD2y
    • Top controller, from left to right: QPD3x, QPD3y, QPD4x, QPD4y
  • tested the remote control and motion of all picomotors

ELOG V3.1.3-